OpenStudy (anonymous):

2. Find 4.84∙39+61∙4.84

OpenStudy (nottim):

use bedmas/pedmas rules. Multiply first, then add.

jhonyy9 (jhonyy9):

this is equal 4,84(39+61)= 4,84 * 100 = 484